A storage tank at STP contains 28.2 kg of nitrogen (N2).

Part A: What is the volume of the tank?

Part B: What is the pressure if an additional 26.6 kg of nitrogen is added without changing the temperature? Please answer in degrees Celsius.
Answer #1

part A)

let the volume of the tank is V

P* V = n * R * T

for STP

1.01 *10^5 * V = 28.2/0.028 * 8.314 * 273

V = 22.6 m^3

the volume of the gas is 22.6 m^3

part B)

if additional mass is added

P * 22.6 = (28.2 + 26.6)/0.028 * 8.314 * 273

solving for P

P = 1.96 *10^5 Pa

the pressure is 1.96 *105 Pa